OpenStudy (anonymous):

Graph f(x)=5x^4+15x^2+10 and use the graph to approximate any zeros (accurate to 3 decimal places) and find zeros algebraically

OpenStudy (ranga):

To find zeros algebraically, put x^2 = t and you will get a quadratic equation in t which you can solve.

OpenStudy (anonymous):

what do you mean? where are you getting the t from?

OpenStudy (ranga):

To find the zeros of f(x)=5x^4+15x^2+10 we need to solve 5x^4+15x^2+10 = 0 This is a fourth degree polynomial and as such it is not easy to solve. But if you recognize this as a quadratic formula in x^2 then we know how to solve a quadratic equation. To make it easy we let t = x^2 and the original equation becomes: 5t^2 + 15t + 10 = 0 (which is a quadratic equation in t). Solve for t first and then put back x^2 in the place of t and solve for x.

OpenStudy (anonymous):

would t = 0?

OpenStudy (ranga):

5t^2 + 15t + 10 = 0 5(t^2 + 3t + 2) = 0 5(t+1)(t+2) = 0 t = -1 or t = -2 put t = x^2 x^2 = -1 and x^2 = -2 Since a square number is negative there are no real solutions for x. So the graph of this function will never cut the x axis because it has no real zeros. But it has complex zeros. +i, -i, +sqrt(2)i and - sqrt(2)i

OpenStudy (anonymous):

thank you @ranga ! but what about the part that says to approzimate any zeros accurate to 3 decimal places??

OpenStudy (ranga):

Since there are no real zeros, there is nothing to be done there.

OpenStudy (anonymous):

oh okay, thank you :)

OpenStudy (ranga):

You are welcome.